# MQuant: Unleashing the Inference Potential of Multimodal Large Language Models via Full Static Quantization

### v1

@@ A.8 Speedup and Memory Savings with Scaling Image Resolution
Table 12: Comparison of perplexity (PPL) with or without online Hadamard transforms in Quarot, evaluated on LLaMA2 models of sizes 7B and 13B. Lower PPL is better.
@@
A.10 Advantage of Proposed MSQ and AIFS
@@
In contrast, per-token dynamic quantization computes quantization parameters on-the-fly for each input token during inference. This approach incurs significantly higher computational overhead, as the quantization parameters must be recalculated for every input token, along with multiple additional memory traversals. Such requirements make per-token dynamic quantization unfriendly or impractical for edge devices and some AI accelerators, which struggle with fine-grained dynamic operations (Tan et al., 2024). This issue is especially severe in MLLMs, where the token count increases significantly with higher image resolution or more video frames. The Modality-Specific Static Quantization (MSQ) in MQuant is a novel per-modality quantization approach specifically designed to address the unique challenges of MLLMs quantization.
@@
A.11 Comparison of Different MLLMs: Input Pre-Process, LayerNorm Architecture in Vision Encoder, Model Parameters and Flops.
Table 13: Comparison of TTFT sensitivity to image resolution, model Parameters and flops in mainstream MLLMs. means the Flops values are measured with the number of visual tokens is 256.
@@
In this section, we compare the visual input pre-process methods, LayerNorm structures in MLLM vision encoder, model parameters, and Flops across five mainstream MLLMs: InternVL2-8B (Chen et al., 2024a), Qwen-VL-Chat-9.6B (Bai et al., 2023b), MiniCPM-V 2.6-8B (Yao et al., 2024), Qwen2-VL-7B (Wang et al., 2024), and GLM-4V-9B (Hong et al., 2024). As shown in Table 13, this comparison highlights the architectural differences between these models, particularly the TTFT sensitivity to input image resolution, illustrating the unique challenges these variations present for quantization and efficient inference.
A.12 Quantization Granularity
Furthermore, as mentioned in SoomthQuant  (Xiao et al., 2022), there are different different granularity levels. The per-tensor static quantization uses a single step size for the entire matrix. Per-token dynamic quantization employs different $s$ for the activations associated with each token, being a common granularity for activations quantization of existing LLMs. For weights, per-channel quantization applies distinct $s$ for each output channel of the weights, while group-wise quantization utilizes a coarse-grained $s$ for different channel groups. Notably, group-wise quantization is a prevalent granularity for weight quantization in LLMs  (Frantar et al., 2022; Yao et al., 2022). Please refer to appendix for more quantization basics.
A.13 LayerNorm, RMSNorm and Computational Invariance
We introduce LayerNorm, RMSNorm, the computational Invariance, and their usage in Transformers.
@@ A.13 LayerNorm, RMSNorm and Computational Invariance
$\text{LayerNorm}(\mathbf{x})=\frac{\mathbf{x}-\mu(\mathbf{x})\mathbf{1}}{\sqrt{\lVert{\mathbf{x}}\lVert_{2}^{2}/d-\mu^{2}(\mathbf{x})+\epsilon}},\text{where}\mu(\mathbf{x})=\frac{\mathbf{1}^{T}\mathbf{x}}{d},\epsilon>0.$ (16)
@@
$\text{RMSNorm}(\mathbf{x})=\frac{\mathbf{x}}{\sqrt{\lVert{\mathbf{x}}\lVert_{2}^{2}/d+\epsilon}},\text{where }\epsilon>0.$ (17)
@@
$\textrm{LayerNorm}(\bm{X})=\textrm{RMSNorm}(\bm{X}\bm{M})\textrm{diag}(\bm{\alpha})\sqrt{D}+\mathbf{1}_{N}\bm{\beta}^{\top}$ (18)
[…]
@@
Based on the computational invariance, recent studies (Ashkboos et al., 2024a, b) have shown that orthogonal transformations can effectively smooth outliers and improve the quantize-ability of both weights and activations. In particular, for transformers, inserting linear layers with an orthogonal matrices $\bm{Q}$ before and after the RMSNorm (Zhang et al., 2019) layer in a transformer, the network remains unchanged. In detail, given the input $\bm{X}$ and orthogonal matrix $\bm{Q}$ for RMSNorm layer, the computational invariance means: $(\bm{X}\bm{Q})\bm{Q}^{\top}=\textrm{RMSNorm}$ RMSNorm. Here, $\bm{Q}^{\top}\bm{Q}=\bm{Q}\bm{Q}^{\top}=\bm{I}$ and a rotation matrix is an orthogonal matrix with $|\bm{Q}|=1$. Note that multiplying a vector $\mathbf{x}$ by $\bm{Q}$ does not change the norm of the vector, since $\|\bm{Q}\mathbf{x}\|=\sqrt{\mathbf{x}^{\top}\bm{Q}^{\top}\bm{Q}\mathbf{x}}=\sqrt{\mathbf{x}^{\top}\mathbf{x}}=\|\mathbf{x}\|$.
A.14 LayerNorm to RMSNorm Transformation.
Figure 8: The proposed Post-LN + Rotate Scheme.
@@
Figure 9: The illustration of transformation from Pre-LN to RMSNorm.

### v2

@@ A.8. Speedup and Memory Savings with Scaling Image Resolution
Table 18. Comparison of perplexity (PPL) with or without online Hadamard transforms in Quarot, evaluated on LLaMA2 models of sizes 7B and 13B. Lower PPL is better.
@@
A.10. Advantage of Proposed MSQ and AIFS
@@
In contrast, per-token dynamic quantization computes quantization parameters on-the-fly for each input token during inference. This approach incurs significantly higher computational overhead, as the quantization parameters must be recalculated for every input token, along with multiple additional memory traversals. Such requirements make per-token dynamic quantization unfriendly or impractical for edge devices and some AI accelerators, which struggle with fine-grained dynamic operations (Tan et al., 2024b). This issue is especially severe in MLLMs, where the token count increases significantly with higher image resolution or more video frames. The Modality-Specific Static Quantization (MSQ) in MQuant is a novel per-modality quantization approach specifically designed to address the unique challenges of MLLMs quantization.
@@
A.11. Comparison of Different MLLMs: Input Pre-Process, LayerNorm Architecture in Vision Encoder, Model Parameters and Flops.
Table 19. Comparison of TTFT sensitivity to image resolution, model Parameters and flops in mainstream MLLMs. means the Flops values are measured with the number of visual tokens is 256.
@@
In this section, we compare the visual input pre-process methods, LayerNorm structures in MLLM vision encoder, model parameters, and Flops across five mainstream MLLMs: InternVL2-8B (Chen et al., 2024b), Qwen-VL-Chat-9.6B (Bai et al., 2023b), MiniCPM-V 2.6-8B (Yao et al., 2024), Qwen2-VL-7B (Wang et al., 2024a), and GLM-4V-9B (Hong et al., 2024). As shown in Table 19, this comparison highlights the architectural differences between these models, particularly the TTFT sensitivity to input image resolution, illustrating the unique challenges these variations present for quantization and efficient inference.
A.12. Quantization Granularity
Furthermore, as mentioned in SoomthQuant  (Xiao et al., 2022), there are different different granularity levels. The per-tensor static quantization uses a single step size for the entire matrix. Per-token dynamic quantization employs different $s$ for the activations associated with each token, being a common granularity for activations quantization of existing LLMs. For weights, per-channel quantization applies distinct $s$ for each output channel of the weights, while group-wise quantization utilizes a coarse-grained $s$ for different channel groups. Notably, group-wise quantization is a prevalent granularity for weight quantization in LLMs  (Frantar et al., 2022).
A.13. LayerNorm, RMSNorm and Computational Invariance
We introduce LayerNorm, RMSNorm, computational Invariance, and their usage in Transformers.
@@ A.13. LayerNorm, RMSNorm and Computational Invariance
(17) $\text{LayerNorm}(\mathbf{x})=\frac{\mathbf{x}-\mu(\mathbf{x})\mathbf{1}}{\sqrt{\lVert{\mathbf{x}}\lVert_{2}^{2}/d-\mu^{2}(\mathbf{x})+\epsilon}},\text{where }\mu(\mathbf{x})=\frac{\mathbf{1}^{T}\mathbf{x}}{d},\epsilon>0.$
@@
(18) $\text{RMSNorm}(\mathbf{x})=\frac{\mathbf{x}}{\sqrt{\lVert{\mathbf{x}}\lVert_{2}^{2}/d+\epsilon}},\text{where }\epsilon>0.$
@@
(19) $\textrm{LayerNorm}(\bm{X})=\textrm{RMSNorm}(\bm{X}\bm{M})\textrm{diag}(\bm{\alpha})\sqrt{D}+\mathbf{1}_{N}\bm{\beta}^{\top}$
where $\textrm{RMSNorm}(\bm{X})$ applies $\mathbf{x}\leftarrow\mathbf{x}/\|\mathbf{x}\|$ to each row of $\bm{X}$, and $\bm{X}=concat(\bm{E}_{v},\bm{E}_{t}))$ is the concatenation between text tokens $\bm{E}_{t}$ and the visual tokens $\bm{E}_{v}$. The vector parameter $\bm{\alpha}$ and offset (vector) parameter $\bm{\beta}$ are learned independently at each LayerNorm instance. The constant matrix $\bm{M}=\bm{I}-\frac{1}{D}\mathbf{1}\mathbf{1}^{\top}$ is a $D\times D$ matrix which subtracts the mean from each row of $\bm{X}$, called recentering operation. Formally, if $\bm{M}=\bm{I}$, the input $\bm{X}$ has a zero-mean, the Eq 19 is equivalent to RMSNorm. Specifically, LayerNorm is widely employed in visual encoders $E$, whereas RMSNorm  (Zhang et al., 2019) is commonly used in LLMs (Touvron et al., 2023a; Dubey et al., 2024) and has been shown to accelerate training and inference time with similar performance (Zhang et al., 2019).
@@
Based on the computational invariance, recent studies (Ashkboos et al., 2024a, b) have shown that orthogonal transformations can effectively smooth outliers and improve the quantize-ability of both weights and activations. In particular, for transformers, inserting linear layers with an orthogonal matrices $\bm{Q}$ before and after the RMSNorm (Zhang et al., 2019) layer in a transformer, the network remains unchanged. In detail, given the input $\bm{X}$ and orthogonal matrix $\bm{Q}$ for RMSNorm layer, the computational invariance means: $\textrm{RMSNorm}(\bm{X}\bm{Q})\bm{Q}^{\top}=\textrm{RMSNorm}$. Here, $\bm{Q}^{\top}\bm{Q}=\bm{Q}\bm{Q}^{\top}=\bm{I}$ and a rotation matrix is an orthogonal matrix with $|\bm{Q}|=1$. Note that multiplying a vector $\mathbf{x}$ by $\bm{Q}$ does not change the norm of the vector, since $\|\bm{Q}\mathbf{x}\|=\sqrt{\mathbf{x}^{\top}\bm{Q}^{\top}\bm{Q}\mathbf{x}}=\sqrt{\mathbf{x}^{\top}\mathbf{x}}=\|\mathbf{x}\|$.
A.14. LayerNorm to RMSNorm Transformation.
Figure 10. The proposed Post-LN + Rotate Scheme.
@@
Figure 11. The illustration of transformation from Pre-LN to RMSNorm.
@@
Figure 12. Illustration the weight distributions for the down-proj layer in different MLLMs’s visual encoder under three conditions: (a) original weight, (b) weights after FHT, and (c) weight after FHT with our RMS.
A.15. Weight Distribution Visualization with RMS.
To facilitate a better analysis of RMS, we also visualize weight distribution across other mainstream MLLMs (InternVL, Qwen-VL, MiniCPM-V and ChatVLM2). As shown in Figure  12, we can observe that the FHT-induced weight outliers presents at all MLLMs. The weight distribution reveals that FHT amplification increases weight magnitudes, while RMS effectively reduces these magnitudes, significantly improving quantization stability.
A.16. Discussion About Video-based MLLMs.
In video-based MLLMs like Video-LLaMA (Zhang et al., 2023) and Qwen2-VL (Wang et al., 2024a), the pipeline involves sampling video frames at fixed intervals, extracting visual features per frame using a vision encoder, and modeling temporal relationships with downstream components (e.g., via temporal positional encoding in Video-LLaMA or 3D convolution in Qwen2-VL). Two key features emerge: (1) videos are processed as multi-frame image sequences, and (2) Our Modality-specific static quantization is not affected by the temporal-aware module. Therefore, for an input video, MQuant’s per-modality static quantization applies naturally: we compute quantization parameters offline across these visual frames. Specifically, the dynamic nature of visual tokens is handled as an image sequence, not altering our static quantization’s efficiency or accuracy in MQuant (Section 3). We recognize this as a promising direction for future work.